OpenStudy (anonymous):

Find two positive even consecutive integers such that the square of the smaller integer is 10 more than the larger integer.

OpenStudy (agl202):

Let x be the smaller x^2 = 10 + (x + 2) x^2 = x + 12 x^2 - x - 12 = 0 (x - 4)(x + 3) = 0 x = 4 or x = -3 The numbers are 4 and 6.